# Epsilon-Greedy Thompson Sampling to Bayesian Optimization

### v2

@@ 2.3 Sampling from Gaussian Process Posteriors
1:input: input variable domain $\mathcal{X}$, number of initial observations $N$, threshold for number of BO iterations $K$, number of spectral points $N_{\text{p}}$, number of sample paths $N_{\text{s}}$, value of $\varepsilon\in(0,1)$
@@
3 Epsilon-greedy Thompson Sampling
In this section, we first introduce another version of TS called sample-average TS (Section 3.1) and then describe $\varepsilon$-greedy TS using the sample-average and generic TS in the context of $\varepsilon$-greedy policy (Section 3.2).
3.1 Sample-average Thompson Sampling
The sample-average TS and the generic TS can be considered two extremes of TS. If $N_{\text{s}}=\infty$, $g({\bf x}|\mathcal{D}^{k-1})$ in Eq. 10 is indeed the GP posterior mean in Eq. 3 whose minimum promotes exploitation. If $N_{s}=1$, the sample-average TS recovers the generic TS that favors exploration. Between these two extremes, therefore, is an optimal state associated with an unknown value of $N_{s}$ at which TS balances exploitation and exploration. However, we do not attempt to find such an optimal state in this work. We instead set $N_{\text{s}}$ at a sufficiently large value, say $N_{\text{s}}=50$, to enforce the exploitation of TS.
3.2 $\varepsilon$-greedy Thompson Sampling
The two distinct extremes of TS motivate the use of $\varepsilon$-greedy policy to randomly switch between them. Specifically, we implement the generic TS with probability $\varepsilon$ to explore the input variable space. We invoke the sample-average TS with probability $1-\varepsilon$ to guide the search toward exploitation. We do not use the posterior mean in Eq. 3 for exploitation because we observe that computing its derivatives is more expensive than computing the derivatives of the average sample path in Eq. 10. It is also straightforward to recover the generic TS from $\varepsilon$-greedy TS by simply setting $\varepsilon=1$ or $N_{\text{s}}=1$. Algorithm 3 details the proposed $\varepsilon$-greedy TS.
(a) Averaging TS, $\varepsilon=0$
(b) $\varepsilon$-greedy TS, $\varepsilon=0.1$
(c) $\varepsilon$-greedy TS, $\varepsilon=0.9$
(d) Generic TS, $\varepsilon=1$
(a) 2d Ackley
(b) 6d Rosenbrock
@@ 4.1 Benchmark Functions
We use the ARD SE covariance function for GP modeling. We call the DACE toolbox [33] to find optimal GP hyperparameters for each minimization problem. We draw $10^{3}$ spectral points to approximate the sample paths of GP posteriors. To minimize the generated sample paths, we use the DIRECT algorithm [45] whose output solution is set as the initial point of an interior-point optimizer. Parameters for the DIRECT algorithm including the function tolerance, maximum number of function evaluations, maximum number of iterations, and maximum number of rectangle divisions are $10^{-9}$, $10^{3}d$, $10^{4}d$, and $10^{4}d$, respectively. Parameters for the interior-point optimizer including the function tolerance, optimality tolerance, maximum number of iterations, maximum number of function evaluations, and termination tolerance on ${\bf x}$ are $10^{-12}$, $10^{-12}$, $200$, $5\times 10^{3}$, and $10^{-12}$, respectively. We carry out all experiments for the benchmark functions using the Carya Cluster of the University of Houston. We compare the optimization results obtained from $\varepsilon$-greedy TS with those from other BO methods, including EI, the lower confidence bound (LCB), the averaging TS, and the generic TS.
To ensure fair comparisons, we randomly generate 100 initial datasets for each problem using Latin hypercube sampling [46], and start each BO method from each of these datasets. In each BO iteration, we record the best-found value of observation error $\log_{10}(y_{\min}-f^{\star})$ and the corresponding solution vector. Here, $y_{\min}$ and $f^{\star}$ represent the best observation of the objective function found in each iteration and its true minimum value, respectively. We set the number of initial observations and the number of BO iterations at $N=5d$ and $K=50$ for the 2d Ackley function, and $N=10d$ and $K=200$ for the 6d Rosenbrock function.
@@
While the optimization results in Fig. 3 suggest that it is safe to set $\varepsilon=0.5$ to balance exploitation and exploration, an optimal $\varepsilon$ may depend on the nature of the problem of interest. A problem that requires more exploitation for its optimal solution (e.g., the 2d Ackley function) may benefit from a smaller $\varepsilon$. A problem with more exploration (e.g., the 6d Rosenbrock function) might be better addressed with a larger $\varepsilon$. Yet determining whether a black-box objective function leans toward exploitation or exploration for its extrema is often elusive. This, ironically, underscores the rationale behind employing $\varepsilon$-greedy TS to prevent our search from overly favoring either exploitation or exploration. It is also worth noting that any endeavors to find an optimal value of $\varepsilon$ for a particular problem should correspond to a specific accuracy level of the initial GP model because lowering the model fidelity always encourages exploration, which is irrespective of the selection of $\varepsilon$.
@@
We also investigate the effect of varying $\varepsilon$ on the runtime for selecting a new solution point. Figure 5 shows approximate distributions of runtime for selecting a new solution point by the sample-average TS with $N_{\text{s}}=50$, $\varepsilon$-greedy TS with different $\varepsilon$ values and $N_{\text{s}}=50$, and the generic TS for the 2d Ackley and 6d Rosenbrock functions. There are two distinct clusters of the runtime values for $\varepsilon$-greedy TS. One cluster aligns with the runtime of the generic TS and the other corresponds to the runtime of the sample-average TS. In addition, the average runtimes for each BO trial of the sample-average TS, $\varepsilon$-greedy TS with $\varepsilon=$ 0.1, 0.5, 0.9, and the generic TS are 124, 102, 64, 18, and 7 s for the 2d Ackley function, respectively. Those for the 6d Rosenbrock functions are 1377, 1188, 809, 229, and 111 s, respectively. We see that, increasing $\varepsilon$ reduces the runtime of each BO trial. This can be attributed to the fact that a larger $\varepsilon$ requires more exploration of calling the generic TS, which is cheaper than the sample-average TS.
@@ 4.2 Cantilever Beam
In this last example, we use $\varepsilon$-greedy TS to identify the constitutive parameters for simulating the behavior of a steel cantilever beam under a static cyclic loading history RH1, as shown in Fig. 6. The web and flange of the beam have different material parameters. In the experiment conducted by Yamada and Jiao [47], forced vertical displacement was applied at the free end of the beam. The defection angle $\theta$ was defined as the ratio of the vertical tip displacement $\Delta$ mm to the beam length $L=800$ mm, i.e., $\theta=\Delta/L$. Values of moment reaction $M$ about $x$-axis at the beam support were measured.
Let ${\bf x}$ denote the parameters underlying a constitutive model for the cyclic behavior of the beam, for which the detailed description is given in Section 4.2.1. Let $M_{t}^{\text{s}}$ denote the simulated value of the moment reaction at the $t$th time step of a discretized loading history of RH1 consisting of $T$ steps, and $M_{t}^{\text{m}}$ the corresponding measured value. We formulate the following misfit function $f(\textbf{x})$ for finding an optimal set of ${\bf x}$ [48]:
@@
To evaluate $f({\bf x})$, we model the beam using Abaqus 2022. We reduce the sensitivity of simulation estimates to the finite-element mesh density by generating a fine mesh of 4960 nodes and 3510 linear hexahedral elements of type C3D8, as shown in Fig. 6. The maximum increment size for each loading history is set as 0.01 s.
@@
4.2.1 Nonlinear Combined Isotropic/Kinematic Hardening
$F=\|\boldsymbol{\xi}\|-\sqrt{\frac{2}{3}}\sigma_{\text{y}}\leq 0,$ (14)
[…]
The isotropic hardening increases the size of the yield surface $F=0$ during the evolution of plastic deformation, while fixing the shape and location of this surface. In this case, $\boldsymbol{\xi}$ in Eq. 14 does not involve the back-stress tensor $\boldsymbol{\alpha}$, resulting in an isotropic yield surface of the stress. This explains why the isotropic hardening cannot capture the Bauschinger effect, which describes an observation that the elastic region associated with monotonically increasing loading is much larger than that associated with reversed loading. Since structural steels exhibit a saturation point of the stress at large deformation, the isotropic hardening for them models the increment of yield surface size using the following monotonically increasing nonlinear function [50]:
where $\sigma_{\text{y},0}$ is the initial yield stress, $Q_{\infty}$ the difference of the stress saturation and $\sigma_{\text{y},0}$, $b$ the isotropic saturation rate, and $\epsilon_{\text{eq}}^{\text{p}}$ the current equivalent plastic strain determined based on its previous state and the rate $\dot{\epsilon}_{\text{eq}}^{\text{p}}$.
Unlike the isotropic hardening, the kinematic hardening does not change the size and shape of the yield surface during the evolution of plastic deformation. Instead, it relocates the center of the yield surface by performing a rigid translation in the evolution direction of the plastic strain. This allows the model to capture the Bauschinger effect. In pure kinematic hardening, the yield function in Eq. 14 has $\sigma_{\text{y}}$ fixed at $\sigma_{\text{y},0}$ and the back-stress tensor $\boldsymbol{\alpha}$ defined as the superposition of $n_{\text{k}}$ back-stress components, such that [51]
Here, the evolution of the $k$th component $\boldsymbol{\alpha}_{k}$ is described by a nonlinear kinematic hardening rule as [52]
Figure 7: Performance of EI, LCB, averaging TS, generic TS, and $\varepsilon$-greedy TS methods for the cantilever beam. (a) Optimization histories; (b) Medians and interquartile ranges of final solutions from 10 runs of each BO method.
Figure 8: Measured and simulated $M$$\theta$ curves of the cantilever beam under different cyclic loading histories. The simulated $M$$\theta$ curves correspond to the best parameter set identified from $\epsilon$-TS and the experimental results from RH1. (a) RH1; (b) RH2; (c) RH3. Here, experimental data obtained from RH2 and RH3 are unseen by the identification process.
@@ 4.2.2 Constitutive Parameters by Bayesian Optimization
During the minimization of misfit function $f(\mathbf{x})$, we keep Young’s modulus and Poisson’s ratio for the web and flange of the beam as constants, set at $E=175.05$ GPa and $\nu=0.3$ [47]. As a result, we identify ten parameters for the beam in which five parameters are associated with the web and the remaining five with the flange. The permissible range for each parameter is taken from [43] and listed in Table 2. To initiate BO, we generate ten distinct initial datasets, half containing 50 data points each and the remainder 100 data points each. With a total of $100$ BO iterations planned, we execute the optimization process using the Carya Cluster housed at the University of Houston. This cluster hosts 9984 Intel CPU cores and 327680 Nvidia GPU cores integrated within 188 compute and 20 GPU nodes. This infrastructure allows us to complete a costly beam simulation in three minutes.
@@
We use the best set of parameters identified from $\epsilon$-greedy TS to predict the $M$$\theta$ curves of the beam under RH1 and other two cyclic loading histories RH2 and RH3. The experimental data associated with RH2 and RH3 were not fed to the parameter identification process. The agreement observed between the measured and simulated $M$$\theta$ curves for each loading history, as shown in Fig. 8, confirms the reliable prediction performance achieved with the identified parameters.
@@ 5 Conclusions
We introduced $\varepsilon$-greedy Thompson sampling (TS) to optimizing costly simulation-based objective functions. The method addresses the classic exploitation–exploration dilemma by randomly switching between its two extremes, namely the generic TS and the sample-average TS. Our empirical findings reveal that $\varepsilon$-greedy TS with an appropriate $\varepsilon$ value robustly achieves the better performance of its two extremes across various problems, and can provide the best result in practical engineering applications.
While several $\varepsilon$-greedy algorithms and the generic TS are guaranteed to converge eventually [28, 4], we look forward to a theoretical analysis to elucidate the convergence properties of $\varepsilon$-greedy TS. Additionally, we are keen on exploring its extensions to high-dimensional settings with support, for example, from a subspace-based approach [53, 54] or a trust-region method [55], and investigating its adaptability to varying $\varepsilon$ during the optimization process.
@@ Acknowledgments
The authors thank the University of Houston for providing startup fund to support this research. Thanks to Prof. Satoshi Yamada at the University of Tokyo and Prof. Makoto Ohsaki at Kyoto University for providing experimental results of the cantilever beam used in Section 4.2.

### v3

@@ 3.2 $\varepsilon$-greedy Thompson Sampling
1:input: input variable domain $\mathcal{X}$, number of initial observations $N$, threshold for number of BO iterations $K$, number of spectral points $N_{\text{p}}$, number of sample paths $N_{\text{s}}$, value of $\varepsilon\in(0,1)$, standard deviation of observation noise $\sigma_{\text{n}}$
@@ 4.1 Benchmark Functions
While $\varepsilon$-greedy TS can be applied to any stationary covariance function listed in Table 1, we restrict our experiments to the ARD SE covariance function, assuming no specific prior knowledge about the characteristics of the objective functions. We call the DACE toolbox [35] to find optimal GP hyperparameters for each minimization problem. We set $\sigma_{\text{n}}=10^{-3}$ for $z-$score standardized output observations and generate $10^{3}$ spectral points to approximate the sample paths of GP posteriors. To minimize the generated sample paths, we use the DIRECT algorithm [48] whose output solution is set as the initial point of an interior-point optimizer. Parameters for the DIRECT algorithm including the function tolerance, maximum number of function evaluations, maximum number of iterations, and maximum number of rectangle divisions are $10^{-9}$, $10^{3}d$, $10^{4}d$, and $10^{4}d$, respectively. Parameters for the interior-point optimizer including the function tolerance, optimality tolerance, maximum number of iterations, maximum number of function evaluations, and termination tolerance on ${\bf x}$ are $10^{-12}$, $10^{-12}$, $200$, $5\times 10^{3}$, and $10^{-12}$, respectively. We carry out all experiments for the benchmark functions using the Carya Cluster of the University of Houston. We compare the optimization results obtained from $\varepsilon$-greedy TS with those from other BO methods, including EI, the lower confidence bound (LCB), the averaging TS, and the generic TS.
To determine an appropriate value of $N_{\text{s}}$ in Eq. 10 which trade-offs between the exploitation and computational cost, our preliminary experiments investigate how changes in $N_{\text{s}}$ affect the optimization results for a fixed value of $\varepsilon$. We observe that when the number of sample paths reaches a sufficiently large value of $N_{\text{s}}=50$, the solutions exhibit less sensitivity to changes in $N_{\text{s}}$.
To ensure fair comparisons, we randomly generate 100 initial datasets for each problem using Latin hypercube sampling [49], and start each BO method from each of these datasets. In each BO iteration, we record the best-found value of observation error $\log(y_{\min}-f^{\star})$ and the corresponding solution vector. Here $y_{\min}$ and $f^{\star}$ represent the best observation of the objective function found in each iteration and its true minimum value, respectively. We set the number of initial observations and the number of BO iterations at $N=5d$ and $K=50$ for the 2d Ackley function, and $N=10d$ and $K=200$ for the 6d Rosenbrock function.
@@
(a) Averaging TS, $\varepsilon=0$
(b) $\varepsilon$-greedy TS, $\varepsilon=0.1$
(c) $\varepsilon$-greedy TS, $\varepsilon=0.9$
(d) Generic TS, $\varepsilon=1$
While the optimization results in Fig. 3 suggest that it is safe to set $\varepsilon=0.5$ to balance exploitation and exploration, an optimal $\varepsilon$ may depend on the nature of the problem of interest. A problem that requires more exploitation for its optimal solution (e.g., the 2d Ackley function) may benefit from a smaller $\varepsilon$. A problem with more exploration (e.g., the 6d Rosenbrock function) might be better addressed with a larger $\varepsilon$. Yet determining whether a black-box objective function leans toward exploitation or exploration for its extrema is often elusive. This, ironically, underscores the rationale behind employing $\varepsilon$-greedy TS to prevent our search from overly favoring either exploitation or exploration. In $\varepsilon$-TS for multi-armed bandits, $\varepsilon$ can be set as $\varepsilon=1/N_{a}$ for achieving simultaneous minimax and asymptotic optimality [31], where $N_{a}$ is the number of arms to select. This, however, does not apply to BO because the number of new solution points, which are similar to arms in multi-armed bandits, is uncountable. It is also worth noting that any endeavors to find an optimal value of $\varepsilon$ for a particular problem should correspond to a specific accuracy level of the initial GP model because lowering the model fidelity always encourages exploration, which is irrespective of the selection of $\varepsilon$.
@@
We also investigate the effect of varying $\varepsilon$ on the runtime for selecting a new solution point. Figure 5 shows approximate distributions of runtime for selecting a new solution point by the sample-average TS with $N_{\text{s}}=50$, $\varepsilon$-greedy TS with different $\varepsilon$ values and $N_{\text{s}}=50$, and the generic TS for the 2d Ackley and 6d Rosenbrock functions. There are two distinct clusters of the runtime values for $\varepsilon$-greedy TS. One cluster aligns with the runtime of the generic TS and the other corresponds to the runtime of the sample-average TS. In addition, the average runtimes for each BO trial of the sample-average TS, $\varepsilon$-greedy TS with $\varepsilon=$ 0.1, 0.5, 0.9, and the generic TS are 124, 102, 64, 18, and 7 s for the 2d Ackley function, respectively. Those for the 6d Rosenbrock functions are 1377, 1188, 809, 229, and 111 s, respectively. We see that increasing $\varepsilon$ reduces the runtime of each BO trial. This can be attributed to the fact that a larger $\varepsilon$ requires more exploration of calling the generic TS, which is cheaper than the sample-average TS.
(a) 2d Ackley
(b) 6d Rosenbrock
@@ 4.2 Cantilever Beam
In this example, we use $\varepsilon$-greedy TS to identify the constitutive parameters for simulating the behavior of a steel cantilever beam under a static cyclic loading history RH1, as shown in Fig. 6. The web and flange of the beam have different material parameters. In the experiment conducted by Yamada and Jiao [50], forced vertical displacement was applied at the free end of the beam. The defection angle $\theta$ was defined as the ratio of the vertical tip displacement $\Delta$ mm to the beam length $L=800$ mm, i.e., $\theta=\Delta/L$. Values of moment reaction $M$ about $x$-axis at the beam support were measured.
Let ${\bf x}$ represent the vector of parameters underlying a constitutive model for the cyclic behavior of the beam, for which the detailed description is given in Section 4.2.1. Let $M_{t}^{\text{s}}$ represent the simulated value of the moment reaction at the $t$th time step of a discretized loading history of RH1 consisting of $T$ steps, and $M_{t}^{\text{m}}$ the corresponding measured value. We formulate the following misfit function $f(\textbf{x})$ for finding an optimal set of ${\bf x}$ [51]:
@@
To evaluate $f({\bf x})$, we model the beam using Abaqus 2022. We reduce the sensitivity of simulation estimates to the finite-element mesh density by generating a fine mesh of 4960 nodes and 3510 linear hexahedral elements of type C3D8, see Fig. 6. The maximum increment size for each loading history is set as 0.01 s.
4.2.1 Nonlinear Combined Isotropic/Kinematic Hardening
$F=\|\boldsymbol{\xi}\|-\sqrt{\frac{2}{3}}\sigma_{\text{y}}\leq 0,$ (14)
[…]
The isotropic hardening increases the size of the yield surface $F=0$ during the evolution of plastic deformation, while fixing the shape and location of this surface. In this case, $\boldsymbol{\xi}$ in Eq. 14 does not involve the back-stress tensor $\boldsymbol{\alpha}$, resulting in an isotropic yield surface of the stress. Since structural steels exhibit a saturation point of the stress at large deformation, the isotropic hardening for them models the increment of yield surface size using the following monotonically increasing nonlinear function [54]:
where $\sigma_{\text{y},0}$ is the initial yield stress, $Q_{\infty}$ the difference of the stress saturation and $\sigma_{\text{y},0}$, $b$ the isotropic saturation rate, and $\epsilon_{\text{eq}}^{\text{p}}$ the current equivalent plastic strain determined based on its previous state and the associated plastic strain rate $\dot{\epsilon}_{\text{eq}}^{\text{p}}$.
Unlike the isotropic hardening, the kinematic hardening does not change the size and shape of the yield surface during the evolution of plastic deformation. Instead, it relocates the center of the yield surface by a rigid translation in the evolution direction of the plastic strain. This allows the model to capture the Bauschinger effect. In pure kinematic hardening, the yield function in Eq. 14 has $\sigma_{\text{y}}$ fixed at $\sigma_{\text{y},0}$ and the back-stress tensor $\boldsymbol{\alpha}$ defined as the superposition of $n_{\text{k}}$ back-stress components, such that [55]
Here the evolution of the $k$th component $\boldsymbol{\alpha}_{k}$ is described by a nonlinear kinematic hardening rule, as [56]
@@ 4.2.2 Constitutive Parameters by Bayesian Optimization
During the minimization of misfit function $f(\mathbf{x})$, we keep Young’s modulus and Poisson’s ratio for the web and flange of the beam as constants, set at $E=175.05$ GPa and $\nu=0.3$ [50]. As a result, we identify ten parameters for the beam of which five parameters are associated with the web and the remaining five with the flange. The permissible range for each parameter is taken from [52] and listed in Table 2. To initiate BO, we generate ten distinct initial datasets, half containing 50 data points each and the remainder 100 data points each. With a total of $100$ BO iterations planned, we execute the optimization process using the Carya Cluster housed at the University of Houston. This cluster hosts 9984 Intel CPU cores and 327680 Nvidia GPU cores integrated within 188 compute and 20 GPU nodes. This infrastructure allows us to complete a costly beam simulation in three minutes.
@@
We use the best set of parameters identified from $\varepsilon$-greedy TS to predict the $M$$\theta$ curves of the beam under RH1 and other two cyclic loading histories RH2 and RH3. The experimental data associated with RH2 and RH3 were not fed to the parameter identification process. The agreement observed between the measured and simulated $M$$\theta$ curves for each loading history, as shown in Fig. 8, confirms the reliable prediction performance achieved with the identified parameters.
Figure 7: Performance of EI, LCB, averaging TS, generic TS, and $\varepsilon$-greedy TS methods for the cantilever beam. (a) Optimization histories; (b) Medians and interquartile ranges of final solutions from ten runs of each BO method.
Figure 8: Measured and simulated $M$$\theta$ curves of the cantilever beam under different cyclic loading histories. The simulated $M$$\theta$ curves correspond to the best parameter set identified from $\varepsilon$-TS and the experimental results from RH1. (a) RH1; (b) RH2; (c) RH3. Here experimental data obtained from RH2 and RH3 are unseen by the identification process.
@@ 5 Conclusions
We introduced $\varepsilon$-greedy Thompson sampling (TS) to optimizing costly simulation-based objective functions. The method addresses the classic exploitation–exploration dilemma by randomly switching between its two extremes, namely the generic TS and the sample-average TS. Our empirical findings reveal that $\varepsilon$-greedy TS with an appropriate $\varepsilon$ value robustly achieves the better performance of its two extremes across the benchmark problems, and can provide the best result in a practical engineering application.
While several $\varepsilon$-greedy algorithms and the generic TS are guaranteed to converge eventually [30, 4], we look forward to a theoretical analysis to elucidate the convergence properties of $\varepsilon$-greedy TS. The regret bound analysis of $\varepsilon$-TS [31] for multi-armed bandits can serve as a starting point for this endeavor. We are also interested in exploring extensions of $\varepsilon$-greedy TS to high-dimensional settings where good approximations of the GP posteriors and efficient optimization of TS acquisition functions in high dimensions are of interest. For this, subspace-based approaches [57, 58] or trust-region methods [59] present viable strategies. Additionally, the application of $\varepsilon$-greedy TS to non-Gaussian likelihood settings and the evolution of $\varepsilon$ during the optimization process are still open problems. The former may be addressed using a Markov chain Monte Carlo algorithm (see e.g., [60, 61]) to sample a large set of function values or to sample the weight values $\boldsymbol{\beta}$ in Eq. 8 from a non-Gaussian posterior.
@@ Acknowledgments
We thank the University of Houston for providing startup fund to support this research. We thank Prof. Satoshi Yamada at the University of Tokyo and Prof. Makoto Ohsaki at Kyoto University for providing experimental results of the cantilever beam used in Section 4.2.